### v2

# CP-Bench: Evaluating Large Language Models for Constraint Modelling

Abstract
Constraint Programming (CP) is widely used to solve combinatorial problems, but its core process, namely constraint modelling, requires significant expertise and is considered to be a bottleneck for wider adoption. Aiming to alleviate this bottleneck, recent studies have explored using Large Language Models (LLMs) to transform combinatorial problem descriptions into executable constraint models. However, the existing evaluation datasets for constraint modelling are often limited to small, homogeneous, or domain-specific instances, which do not capture the diversity of real-world scenarios. This work addresses this gap by introducing CP-Bench, a novel benchmark that includes a diverse set of well-known combinatorial problems sourced from the CP community, structured explicitly for evaluating LLM-driven CP modelling. With this dataset, and given the variety of constraint modelling frameworks, we compare and evaluate the modelling capabilities of LLMs for three distinct constraint modelling systems, which vary in abstraction level and underlying syntax. Notably, the results show higher performance when modelling with a high-level Python-based framework. Additionally, we systematically evaluate the use of prompt-based and inference-time compute methods across different LLMs, which further increase accuracy, reaching up to 70% on this highly challenging benchmark.
1 Introduction
Combinatorial problems exist in numerous real-world applications, where optimal decision-making is essential, including logistics, scheduling, and network design [29]. Solving these problems requires finding the best combination of elements under specific constraints, which is computationally intensive; many such problems are NP-hard, meaning that their solving time can increase exponentially with problem size in the worst case.
Constraint Programming (CP) is an established paradigm for solving combinatorial problems, employing a declarative approach in which users specify what constraints the solution must satisfy, rather than detailing how to find it [40, 35]. Despite its solving effectiveness, the complex process of modelling—translating a problem description into a formal CP model—remains a significant bottleneck. It involves identifying and formally defining the problem’s decision variables with their domains, constraints on the variables, and, where applicable, an objective function. This process demands a deep understanding of the application domain and proficiency in the chosen modelling framework, both syntactically and semantically. This modelling complexity is recognised as one of the most important limitations for the wider use of combinatorial optimisation techniques, restricting accessibility for non-experts [11, 10].
Figure 1: LLM-driven constraint modelling: users state a combinatorial problem in natural language, which the system transforms into a formal constraint representation, and delegates the latter to a constraint solver.
This raises a central question: Can we lower the modelling barrier to make CP accessible to a broader audience? Prior works in this direction include constraint acquisition methods that aim to learn the constraints from data [2, 3, 23] and constraint detection methods from natural language [18]. More recently, advances in large language models (LLMs) offer the opportunity of creating modelling assistants that could initiate or (semi-)automate the modelling process (as illustrated in Figure 1), in analogy to how coding assistants are widely being used for a variety of coding tasks [17, 6, 9].
Nevertheless, LLM-driven constraint modelling presents different challenges compared to general code generation. First, while imperative code specifies step-by-step sequential instructions, declarative modelling requires that all constraints must hold true simultaneously. Second, formalizing the extracted CP components is hard due to the combinatorial nature of modelling choices themselves. Specifically, modelling requires selecting a viewpoint—a specific way to represent variables and structure constraints based on the problem domain—from potentially many alternatives [12]. Third, verifying the correctness of a generated constraint model is different from imperative code generation, in which correctness can often be tested using example inputs and expected outputs (e.g., unit tests). In contrast, constraint modelling lacks such verifications, as a “test” would require knowing the correct model or solution in advance.
Recent works have explored the potential of LLMs in this direction by generating solvable representations of numerical reasoning problems in formal declarative modelling frameworks [43, 38]. Ishay et al. [15] frame Logic Grid Puzzles (LGPs) as Answer Set Programming (ASP) programs, Prasath and Karande [32] focus on linear optimization problems, and Ye et al. [43] propose translating inputs into satisfiability problems. Although promising results have been achieved for simple domain-specific problems, such as small-scale linear optimization problems [34, 32] and LGPs [15], scaling to more complex and diverse problems remains a significant challenge [24].
Existing limitations can be summarized as follows. (a) Current NL-to-CP benchmarks lack the diversity and complexity necessary to represent practical combinatorial problems. This constitutes a prerequisite for assessing LLM capabilities on this task. (b) Given the wide variety of constraint modelling frameworks available, ranging in abstraction levels and interface types, there is a need for systematic evaluation across them to explore their impact on LLMs. (c) Applying recent LLM advances to combinatorial modelling remains underexplored, even though test-time scaling methods have shown success for complex programming tasks [7].
To address the lack of representative benchmarks, we created CP-Bench, a novel dataset of diverse combinatorial problems gathered from distinct sources within the CP community. While existing repositories (e.g., PyCSP3 examples [20], CSPLib [13], and others) contain models, they are not structured for automated evaluation of LLM-generated models. We designed CP-Bench for this task based on solution-level evaluation, which is non-trivial since more than one (optimal) solution may exist for a combinatorial problem (detailed in Section 5.2).
Using this benchmark, we systematically evaluate the capabilities of state-of-the-art LLMs when prompted to generate models using different frameworks, varying by abstraction level (low vs. high) and interface type (domain-specific vs. Python-based). We select the following as representatives: OR-Tools [31, 30] offers a low-level, direct solver Python API; CPMpy [14] provides a high-level, Python-based modelling interface; and MiniZinc [27] is a well-established high-level, domain-specific modelling language. Across these frameworks, we compare different system prompts, starting from basic minimal instructions and progressing to detailed guidelines and including documentation specific to the framework used.
With inference-time computation showing high success rates across various LLM tasks [37], we adapt and evaluate such methods in order to further enhance the quality of the generated constraint models. Firstly, we use Retrieval-Augmented In-Context Learning (RAICL) to enrich the prompt context with representative examples [44, 24]. Secondly, we exploit the probabilistic nature of LLMs, by repeatedly sampling multiple CP models and employing solution majority voting to select the most reliable model based on the observed outputs [6, 28]. Finally, to address detectable errors, such as runtime or solution printing ones, and trigger self-correction of modelling errors, we investigate the effect of self-verification prompting on LLMs for iterative model refinement.
The primary contributions of this work are the following:
  • A novel benchmark comprising 101 diverse combinatorial problem descriptions along with their runnable CP models, structured specifically for evaluating LLM-driven constraint modelling.
  • Systematic evaluation of state-of-the-art LLMs across three commonly used modelling frameworks, showing the impact of interface type on accuracy, with Python-based frameworks achieving up to 65% accuracy compared to a maximum of 50% for the domain-specific language.
  • Adaptation of prompt-based and inference-time compute methods for enhancing LLM-driven constraint modelling, reaching up to 70% accuracy.
2 Background
2.1 Constraint Programming
We use CP as a formal intermediate representation for the given combinatorial problems. A Constraint Optimization Problem (COP) is a type of combinatorial problem that involves finding an optimal assignment of values to decision variables, subject to a set of constraints and an objective function. Formally, a COP is defined as a tuple $(\mathbf{X},\mathbf{D},\mathbf{C},f)$ where:
  • $\mathbf{X}=\{x_{1},x_{2},\ldots,x_{n}\}$ is a set of $n$ decision variables.
  • $\mathbf{D}=\{D_{1},D_{2},\ldots,D_{n}\}$ is a set of $n$ domains, where each $D_{i}$ is a finite set of allowable values for $x_{i}$.
  • $\mathbf{C}=\{C_{1},C_{2},\ldots,C_{m}\}$ is a set of $m$ constraints, with each $C_{j}$ specifying allowed combinations of values for a subset of variables. Formally, $C_{j}\subseteq D_{j1}\times D_{j2}\times\ldots\times D_{jk}$ for some subset $\{x_{j1},x_{j2},\ldots,x_{jk}\}\subseteq\mathbf{X}$.
  • $f:\prod_{i=1}^{n}D_{i}\rightarrow\mathbb{R}$ is the objective function to be optimized (either maximized or minimized).
The goal is to find an assignment $\mathbf{a}=\{x_{1}=v_{1},x_{2}=v_{2},\ldots,x_{n}=v_{n}\}$ that satisfies all constraints in $\mathbf{C}$ and optimizes the objective function $f$. Formally, an assignment $\mathbf{a}$ is feasible if $v_{i}\in D_{i}$ for all $i$ and if $(v_{j1},v_{j2},\ldots,v_{jk})\in C_{j}$ for each constraint $C_{j}$. An optimal solution $\mathbf{a}^{*}$ is a feasible assignment that optimizes $f$. If no objective function $f$ is defined, the problem reduces to a Constraint Satisfaction Problem (CSP), where the goal is to find any feasible assignment. When no feasible assignment exists, the problem is considered unsatisfiable. For brevity, we will use $\mathbf{a}^{*}$ to represent either an optimal solution in COPs or a feasible solution in CSPs, and we will commonly refer to it as the optimal solution.
Notably, CP systems are not restricted to linear constraints but also support a variety of global constraints that capture common patterns in combinatorial problems. These include min/max aggregators, element constraints (for array access), count/nr-values (for aggregation over arrays), cumulative constraints (for scheduling-type problems), and many others that facilitate more expressive modelling of complex problems [1].
2.2 Large Language Models
Based on the Transformer architecture [39], LLMs are deep learning systems, typically with billions of parameters, capable of learning and generating complex language patterns [21]. In this work, we use them as black-box text generators, configured only by an input text (or prompt) and a randomness parameter (temperature). Formally:
$\begin{split}LLM(p,\tau)=(x_{1},x_{2},\ldots,x_{w}),\quad\text{where }\\ x_{t+1}\sim P_{\tau}(x_{t+1}\mid p,x_{\leq t})\end{split}$ (1)
Here, $x_{\leq t}$ denotes the sequence of tokens generated up to time step $t$, with $x_{\leq 0}$ as the empty sequence and $x_{1}$ as the first generated token. At each step, the next token $x_{t+1}$ is sampled from the temperature-adjusted probability distribution $P_{\tau}(x_{t+1}\mid p,x_{\leq t})$, which is typically computed by applying the softmax function to the model’s logits—raw unnormalized scores over the vocabulary. The temperature $\tau$ scales the logits before softmax, with higher values increasing randomness and diversity, while lower values favour the most likely tokens. The generation process continues until a predefined stop token is encountered or the maximum context length is reached. We assume tokenization and detokenization into text are handled internally by the LLM.
3 Problem Formulation
As illustrated in Figure 1, we conceptualize a system capable of assisting users in modelling combinatorial problems, transforming their textual descriptions into executable models that a constraint solver can solve. In this paper, we focus on this transformation, which requires a mechanism that can process and accurately represent the problem’s constraints, variables, and objectives—tasks that demand expert knowledge in both the application domain and CP modelling. We consider LLMs such versatile tools, because of their powerful natural language processing and coding capabilities [6].
Formally, we define the process as follows: Let $\mathcal{V}^{*}$ represent the set of all possible text sequences that can be composed using tokens from the LLM’s vocabulary $\mathcal{V}$. Let $\mathcal{M}$ be the space of all possible constraint models, and let $\mathcal{S}$ be the space of all possible valid assignments (solutions) to the decision variables of a constraint model. We utilize the previously defined function $LLM$, which maps an input prompt [^1] We assume that input $p$ includes a problem description, parameter values (for parameterized problems) and instructions for a modelling framework. $p\in\mathcal{V}^{*}$ and a temperature parameter $\tau\in\mathbb{R}$ to an output text that contains an instantiated constraint model $m\in\mathcal{M}$. Next, we define a function $Solver:\mathcal{M}\rightarrow\mathcal{S}\cup\{\emptyset,error\}$. Given an instantiated constraint model $m$, the Solver yields an output $\mathbf{a}^{+}$, which represents either an optimal solution $\mathbf{a}^{*}\in\mathcal{S}$, an unsatisfiable model $\emptyset$, or some kind of solver error (e.g., syntax or runtime). The overall system can be represented as follows:
$Solver(LLM(p,\tau))=\mathbf{a}^{+}$ (2)
4 Methodology
The task for LLMs in this context is to translate problem descriptions into a declarative specification of variables and constraints over them, including expressing them correctly in a specific modelling framework and generating code for printing the computed solution in a structured way. These latter requirements demand coding skills; thus, we take inspiration from successful test-time scaling techniques used for generating imperative code.
4.1 System Prompt
Given the complexity of translating natural language descriptions into constraint models, we explore different levels of system prompts to guide LLMs more effectively in a 0-shot setting. Without examples in the prompt (hence 0-shot), providing a comprehensive description of the task is important to help the LLM process and correctly model the problem declaratively. Formally, if $sys\in\mathcal{V}^{*}$ is the system prompt describing the task, and $c\in\mathcal{V}^{*}$ is the combinatorial problem description, then the LLM input prompt is $p=sys\oplus c$, concatenating the system prompt and the problem description. We examine three levels of system prompts, each progressively providing more detailed information (the prompts are available in the supplementary material [25]):
  1. 1.
    Basic prompt. The basic prompt provides essential instructions to the LLM, only describing the task of producing a constraint model using a specific framework and how the solution should be formatted. This prompt simulates a scenario where a user gives minimal guidance, e.g. as one would do in a chat interface.
  2. 2.
    Guidelines prompt. Expanding on the basic prompt, this level adds specific guidelines related to the modelling process. These guidelines include code generation steps, generic classroom-style advice on constraint modelling, and a template on how to format the response model in the modelling framework used. It mirrors a situation where a user provides more elaborate but still generic instructions.
  3. 3.
    Documentation prompt. The most detailed prompt builds on top of the guidelines prompt, appending single-line API documentation of the available classes and functions of the CP modelling framework. This prompt offers (limited) access to solver documentation, enabling direct reference to available methods and functionalities.
4.2 In-Context Learning
Few-shot prompting is a widely used method that includes task-relevant examples in the input prompt, improving LLM responses without requiring retraining [4]. However, selecting effective examples is critical, as a static or random set of examples may not align well with diverse inputs [22, 42]. To address this, recent studies have examined dynamically inserting examples in the prompt context based on the current input [44]. In the context of constraint modelling, adding examples of semantically similar problems and their corresponding models has previously improved the accuracy of generated models [24].
Figure 2: Retrieval-augmented few-shot prompting: on-the-fly example retrieval to provide more relevant patterns to the LLM.
Formally, this method retrieves from a prebuilt database an ordered set $E$ of $e$ input-output pairs, $E=\left\{(i_{j},o_{j})\right\}_{j=1}^{e}$, based on their relation to the given problem $c$. Given a database of problem-model pairs, the retrieval process involves comparing the token embeddings of $c$ with each problem in the database and selecting the $e$ most similar ones using a predefined metric, such as semantic similarity [5]. The prompt then becomes $p=sys\oplus E\oplus c$, where the retrieved examples $E$ are placed between the system prompt and the current problem description, as shown in Figure 2.
4.3 Repeated Sampling
While enriching the prompt with instructions and examples can enhance the guidance provided to LLMs, it does not fully address the complexity of synthesizing CP models from textual descriptions.
Algorithm 1 Sampling with Solution Majority Voting
[…]
2:for $i=1$ to $k$ do
3:  $m_{i}\leftarrow LLM(p,\tau)$  // Generate CP model
4:  $s_{i}\leftarrow Solver(m_{i})$  // Extract the solution
5:  $\mathcal{C},\mathcal{S}\leftarrow\mathcal{C}\cup\{m_{i}\},\mathcal{S}\cup\{s_{i}\}$
6:end for
[…]
8:if $s_{\text{maj}}$ exists then
9:  return  $\mathcal{C}[\text{first\_index\_of}(s_{\text{maj}},\mathcal{S})]$
10:end if
11:return  $\mathcal{C}[1]$  // Fallback to first candidate model
In this direction, we aim to benefit from the probabilistic nature of LLMs through repeated sampling [6]. Instead of relying on the single “most probable” output, we increase the temperature parameter value $\tau$, generating multiple candidate models for the same problem description. This strategy exploits the variability in LLM outputs to identify the most consistent model across multiple runs.
Specifically, we apply Solution Majority Voting for the final selection of the generated model [28]. By repeatedly querying the LLM and solving each generated constraint model, we observe the solutions they produce; then, the first model yielding the most frequently observed solution is selected. If none of the models produces a solution, we default to the first generated model. The process is outlined in Algorithm 1.
4.4 Self-Verification
Writing semantically and syntactically correct code in a single pass, without the ability to revise previous steps, is a complex task even for experienced programmers. This challenge holds even more for CP modelling, where variables need to be declared before constraints, and constraints should not contradict earlier ones. We expect that the generated models may occasionally contain runtime, modelling or formatting errors. This motivates us to define three criteria that assert the validity of the generated constraint model:
  • Runtime Integrity: The generated code executes without errors and correctly utilizes library components.
  • Model Accuracy: The decision variables, constraints, and objective function are correctly defined and aligned with the problem description; when the model is executed, the produced solution is valid and optimal.
  • Solution Output: The solution is printed in the correct structured format as per the user’s provided instructions.
A promising approach to improve the reliability of code generation is to allow LLMs to review, verify, or refine their previous output, effectively mimicking the human debugging process [7, 41].
We composed a Self-Verification base prompt $v$ that guides the LLM through an evaluation and correction process, instructing it to iteratively assess the model based on the aforementioned criteria. Appended to this prompt are: (a) the system prompt syssys, (b) the problem description $c$, (c) the model $m$ generated by the LLM in the previous attempt, and (d) its resulting output $\mathbf{a}^{+}$ when executed. Formally, the full self-verification prompt at iteration $t$ is denoted as $v\oplus sys\oplus c\oplus m_{t-1}\oplus\mathbf{a}^{+}_{t-1}$. This process continues until either the LLM determines that the most recent constraint model $m_{t-1}$ is correct according to the specified criteria, or a predefined number of iterations is reached; at that point, the latest model is selected. The process is depicted in Figure 3.
Figure 3: Iterative Self-Verification: the generated CP model is verified iteratively; both the problem and produced solution are also provided back to the LLM. In this work, we use a single LLM for both model generation and verification (thus, self-verification).
5 Dataset: CP-Bench
Table 1: Comparison of CP-Bench with existing datasets for evaluating LLM-generated CP models of combinatorial problems. Unique constraints count distinct constraint operators and their argument structures (arity and type, e.g., variable/constant/expression).
Benchmark # Constraints Decision Variables # Unique # Optimization
Mean (Median, IQR) Min/Max Mean (Median, IQR) Min/Max Constraints Problems
NL4Opt 289 2.90 (3.0, 1.0) 2/5 2.02 (2.0, 0.0) 2/3 27 All
LGPs 100 38.47 (38.0, 29.0) 7/69 12.00 (12.0, 0.0) 12/12 8 None
CP-Bench [ours] 101 84.76 (13.5, 45.0) 1/2017 51.69 (13.0, 17.3) 3/962 241 30
Table 2: CP-Bench sources. Selection was based on CPMpy model availability to facilitate evaluation (the ground-truth model is necessary).
Source # Selected Problems
CSPLib 25 IDs: 1, 2, 5–13, 15, 19, 21, 24, 26, 28, 33, 44, 49, 50, 53, 54, 76, 84.
CPMpy Examples 16 bus_scheduling, minesweeper, packing_rectangles, jobshop, knapsack, mario, n_puzzle, RCPS, room, send, set_game, sudoku, tsp, agatha, wolf, zebra.
Håkan K. 42 From 3_coins to crypta (alphabetically), excluding test files.
Course 18 All problems from the repository.
The evaluation of LLMs on translating textual problem descriptions into formal CP specifications has so far been limited by the scope and diversity of available datasets. Existing benchmarks, such as NL4Opt [33]–part of a competition at NeurIPS [^2] https://neurips.cc/virtual/2022/competition/50079– and Logic Grid Puzzles (LGPs) [16], primarily feature small, domain-specific problems, such as simple linear programming (LP) problems or homogeneous puzzles, as summarized in Table 1. More recently, Singirikonda et al. [36] presented a cross-domain dataset with MiniZinc models, focusing on CP, LP and Mixed-Integer Programming (MIP) problems as well.
To address the gap in CP-modelling LLM benchmarks [24], we introduce CP-Bench [^3] https://huggingface.co/datasets/kostis-init/CP-Bench. [26], a collection of 101 diverse and challenging combinatorial problems drawn from well-established sources in the CP community, as detailed in Table 2. To achieve diversity, we selected problems from the popular “benchmark library for constraints” CSPLib [13], from the online examples repository of the CPMpy modelling framework [14], from the extensive online repository of Håkan Kjellerstrand [19], and from a modelling course-based set of problems [24].
Table 1 highlights the diversity and complexity of CP-bench compared to existing benchmarks. It spans optimization and satisfaction problems, with a wide range of decision variables, constraints, and 241 distinct constraint types. In particular, the number of constraints per problem ranges from 1 to 2017, and the number of variables from 3 to 962, offering a significantly broader scope than NL4Opt and LGPs (Table 1).
5.1 Instance Structure
Each dataset instance was transformed into a self-contained Python file containing all the necessary components for evaluating the modelling capabilities of LLMs, with the following structure:
  1. 1.
    Metadata: The source(s) of the problem along with any additional relevant information.
  2. 2.
    Problem Description: The original problem statement as provided by the source, with an additional print instruction that specifies the required output structure.
  3. 3.
    Input Data (optional): For problem descriptions that lack parameter instantiation, one valid data instance is provided in plain Python syntax, using only built-in types.
  4. 4.
    Model: The decision variables, constraints, and objectives (if any) that make up a runnable CP model of the problem, using the CPMpy modelling framework.
  5. 5.
    Solution Printing: Solution printing code, following the requirements in the problem description.
Importantly, only the Problem Description and (optionally) Input Data are provided to the LLM, the remaining components are used exclusively for evaluation.
5.2 Evaluation Process
As the focus lies on problems that may have more than one optimal solution, it is mandatory to adopt an appropriate metric to evaluate the generated constraint models. Related works have used constraint-level, model-level, and solution-level accuracy [33, 24]. The first two metrics require mapping the decision variables of the generated model with those from the ground-truth model. However, when modelling complex problems, multiple possible viewpoints and formulations of the decision variables are possible, thus making it highly challenging to map variables between the generated and ground-truth models in a correct but generic way. This leads us to opt for solution accuracy, which measures the correctness of the solution obtained by executing the generated code; correctness is evaluated with respect to the ground-truth model, considering both feasibility and optimality. Additionally, this metric allows for evaluating LLMs on any modelling framework desired, as long as a final solution can be printed.
Formally, let $N$ be the total number of test instances, $\hat{\mathbf{a}}_{i}$ the solution produced by the generated CP model of instance $i$, and $\mathbf{a}_{i}^{*}$ an optimal solution from the ground truth constraint model $\mathbf{C}_{i}$. Solution accuracy is defined as:
$\text{SA}(\hat{\mathbf{a}})=\frac{1}{N}\sum_{i=1}^{N}\mathbb{I}\left(\hat{\mathbf{a}}_{i}\in\text{Sol}(\mathbf{C}_{i})\wedge f_{i}(\hat{\mathbf{a}}_{i})=f_{i}(\mathbf{a}_{i}^{*})\right)$ (3)
Here, the term $\hat{\mathbf{a}}_{i}\in\text{Sol}(\mathbf{C}_{i})$ tests whether the generated solution satisfies all constraints of the ground-truth model. For optimization problems, the second condition $f_{i}(\hat{\mathbf{a}}_{i})=f_{i}(\mathbf{a}_{i}^{*})$ ensures equivalence to the optimal objective value.
6 Experiments
In our experiments, we focus on the following questions:
  1. Q1:
    How do LLMs perform in generating constraint models across different modelling frameworks?
  2. Q2:
    What types of errors do LLMs make when generating constraint models, and how do different system prompts address them?
  3. Q3:
    What is the impact of inference-time compute methods, such as few-shot prompting, repeated sampling, and self-verification, individually and in combination?
6.1 Experimental Setup
Table 3 lists the LLMs used throughout the experiments. We considered covering a range of properties (coding, reasoning, rank, mixture of experts, number of parameters, open weights/proprietary, context window size, etc.). Specifically, we selected 8 models from the LMSys Chatbot Arena leaderboard focused on coding tasks [8]. [^4] As accessed on 04 May 2025. Notably, gpt-4.1-mini was preferred to gpt-4.1 (it performed almost equivalently) and gpt-4.1-nano (it showed very low performance). We used [^5] The code is available here: https://github.com/kostis-init/CP-Bench the OpenAI [^6] https://platform.openai.com/docs/api-reference, Together AI [^7] https://docs.together.ai/reference, Anthropic [^8] https://docs.anthropic.com/en/api, and DeepSeek [^9] https://platform.deepseek.com/docs APIs via their Python client libraries, using Python 3.12. For executing the generated constraint models, we used a 10-second timeout due to the small instance sizes of the problems. Finally, we set the temperature parameter $\tau$ to 0 for deterministic outputs, except when applying repeated sampling techniques, where we increased it to $\tau=0.8$ for variability, following existing works [28].
Table 3: LLMs used in the experiments. Parameter counts and context sizes are approximate, ’*’ denotes active parameters for Mixture-of-Experts models, and N/A means we found no public information about it.
LLM Org. # Params Context Reasoning
gpt-4.1-mini-2025-04-14 OpenAI N/A 1M -
o4-mini-2025-04-16 N/A 200k
DeepSeek-V3-0324 DeepSeek 37B* 64k -
DeepSeek-R1 37B* 64k
Qwen2.5-Coder-32B Alibaba 32B 128k -
QwQ-32B 32B 132k
Llama3.3-70B Meta 70B 128k -
Llama4-Maverick 17B* 1M -
6.2 Results & Discussion
We now present the results and discuss them.
6.2.1 Q1: Modelling Framework
Figure 4: Percentages of successfully generated models. From left to right: MiniZinc, CPMpy, OR-Tools. Reasoning LLMs are prepended with (r).
Figure 5: Average Accuracy by Framework and System Prompt Level (Aggregated across LLMs).
Figure 4 shows the detailed results, and Figure 5 illustrates, aggregated across all LLMs, the average percentage of successfully generated models for each modelling framework (MiniZinc, CPMpy, and the OR-Tools CP-SAT solver) for the different system prompt levels.
The results indicate that the choice of modelling framework influences the success rates of LLMs: when instructed to utilize the CPMpy library for modelling, they exhibit a consistent improvement in performance compared to modelling with MiniZinc and OR-Tools, especially with more detailed system prompts (Levels 2 and 3). The higher accuracy observed with CPMpy is potentially attributable to two factors. First, contemporary LLM training tends to include high quantities of Python code [6]; this familiarity with Python syntax and libraries facilitates generating models in Python-based frameworks. This is also a factor for the increased performance with OR-Tools compared to MiniZinc. Second, CPMpy provides higher-level modelling abstractions compared to OR-Tools, including nested expressions, double reification, as well as a larger set of built-in constraints that can be used. In contrast, the technical low-level interface of OR-Tools often poses challenges for LLMs, leading to reduced accuracy despite the large amount of available documentation and examples.
Looking at the individual performance of LLMs across the frameworks in detail (Figure 4), we can see that the leading models, GPT-4.1-mini and o4-mini, demonstrate strong capabilities, particularly when generating CPMpy or OR-Tools code, peaking at 63% and 65% respectively. DeepSeek-V3, DeepSeek-R1 and QwQ-32B follow similar trends, although with relatively lower accuracy, achieving their best results with CPMpy (all around $\sim$53% at Level 3). These LLMs benefit from increased prompt detail across the high-level frameworks, but not for the low-level OR-Tools API, where Level 2 prompts (modelling guidelines only) were most effective for DeepSeek-V3 and QwQ-32B. The remaining LLMs generally show lower performance across the frameworks.
Focusing on reasoning LLMs, we observe that o4-mini outperforms all other LLMs on average across frameworks, while QwQ-32B, despite its relatively small size, performs on par with models like GPT-4.1-mini and DeepSeek-V3, and even outperforms both Meta’s Llama models. Interestingly, reasoning models show a relative advantage when modelling with low-level or domain-specific frameworks, as DeepSeek-R1 and o4-mini are the only ones to score above 45% for MiniZinc, and 50% for OR-Tools, while for CPMpy, their non-reasoning counterparts achieved higher accuracies.
In summary, generating models using Python libraries (CPMpy and OR-Tools) tends to yield higher accuracy for most LLMs than the domain-specific MiniZinc language. We note that these results reflect the capabilities of current-generation LLMs; ongoing advances in model architecture and training data might improve accuracy in future iterations.
6.2.2 Q2: System Prompt & Error Types
Figures 4 and 5 also demonstrate the impact of varying system prompt levels across frameworks and LLMs. A key observation across all configurations is that providing at least some specific guidance — either classroom-level guidelines (Level 2) or framework documentation (Level 3) — consistently yields higher accuracy than the basic prompt (Level 1). However, the addition of detailed documentation on top of the guidelines showed some variations. For high-level frameworks like MiniZinc and CPMpy, incorporating API documentation in the system prompts led to an increase in accuracy for 6 out of 8 LLMs. This indicates that more documentation is beneficial to produce valid constraint models when using high-level modelling systems. Interestingly, for the lower-level OR-Tools solver interface, while adding guidelines (Level 2) improved performance over the baseline (Level 1) for all LLMs, further adding detailed documentation (Level 3) did not consistently improve upon Level 2. In fact, average performance slightly decreased, with only 3 out of 8 LLMs showing any benefit from the extra documentation on OR-Tools. This indicates that the LLMs might struggle with integrating extensive technical details of lower-level modelling frameworks.
To better understand the impact of prompt levels, we examine the types of errors made under different prompt configurations, depicted in Figure 6. We categorize the types of errors into two main categories: detectable and modelling errors. Detectable errors include basic failures like failing to execute the generated code, problems parsing outputs, or running into timeouts. Modelling errors are those instances where either the generated model is unsatisfiable or it produces a wrong solution (according to our solution-level evaluation metric, Section 5.2). For CPMpy, as the prompt level increases, detectable errors decrease while modelling errors rise. A similar trend was observed for MiniZinc and OR-Tools (figures are available in the Appendix). The reduction in detectable errors with more detailed prompts suggests that LLMs generate more syntactically correct and executable code as they receive more information. This partly leads to more successes, but at times also to an increase in logical modelling errors (bottom subfigure), where the code is syntactically correct but the solution is either incorrect or not found.
Figure 6: Detectable and modelling errors with CPMpy.
6.2.3 Q3: Inference-Time Compute Methods
Finally, we assess the impact of various inference-time strategies from Section 4, using the following configurations:
  1. 1.
    Baseline (Section 4.1): The input consists of the system prompt and problem description. The first generated model is directly evaluated without further processing, as in Q1.
  2. 2.
    RAICL (Section 4.2): The input contains the system prompt and $e=8$ examples based on reverse semantic similarity, following existing work [24]. We use a leave-one-out strategy for evaluation, testing each problem individually while utilizing the 100 remaining ones as the retrieval database.
  3. 3.
    Sampling (Section 4.3): The input consists of the system prompt and problem description. The LLM generates $k=10$ responses, with the final model selected through solution majority voting.
  4. 4.
    Self-Verification (Section 4.4): A model is generated with the Baseline method and the LLM iteratively refines it through self-verification, with a maximum number of 10 iterations.
  5. 5.
    Sampling & Self-Verification: The Sampling method generates a model, which is input to Self-Verification.
For cost reasons, we included four LLMs (one from each organisation): gpt-4.1-mini, DeepSeek-V3, QwQ-32B, Llama-3.3-70B, and we used CPMpy due to its higher accuracy on average (Q1). Based on the Q2 results, we used the documentation system prompt (Level 3). The results are shown in Figure 7.
Surprisingly, the RAICL strategy was shown ineffective for all tested LLMs; performance consistently decreased compared to the Baseline when in-context examples were provided. Based on these results and compared to the results from Michailidis et al. [24], it seems that LLMs handle more effectively the given structured framework documentation in the system prompt than they generalize from specific problem-model examples provided in context. This simplifies practical deployment by removing the need for custom example database construction and retrieval methods. As such, investigating more modelling tools or frameworks could focus on merely adapting their documentation for the system prompt.
Conversely, the other inference-time compute methods boosted performance across all LLMs, with the most significant gains observed by Qwen’s reasoning LLM QwQ-32B (approx. 20% increase). Repeated sampling improved performance across all LLMs, often showing approximately a 10% increase in accuracy. Notably, with this method, the reasoning-trained QwQ-32B performed on par with DeepSeek’s and OpenAI’s larger LLMs. Self-verification also improved accuracy significantly compared to the Baseline, yielding results comparable to repeated sampling, although gpt-4.1-mini performed slightly better with sampling than with self-verification alone. As RAICL negatively affected accuracy, we experimented with the combination of the best-performing methods: repeated sampling & self-verification. This combined approach further increased accuracy, with all LLMs tested demonstrating their highest performance in this setting.
Figure 7: Performance across different inference-time compute methods for each LLM using CPMpy.
Overall, the highest accuracy is achieved when prompting gpt-4.1-mini with the documentation prompt and the combined configuration, resulting in 71/101 correct CP models in this 0-shot setting; only 2 instances resulted in detectable errors (1 runtime error, 1 max tokens exceeded), and 28 instances resulted in modelling errors (unsatisfiable or wrong solution). As this is a rapidly advancing area, we have set up an online leaderboard [^10] https://huggingface.co/spaces/kostis-init/CP-Bench-Leaderboard of the benchmark, on which evaluations with newer LLMs and alternative model generation methods can be submitted.
7 Conclusion
In this work, we introduced a diverse dataset of combinatorial problems and systematically evaluated various state-of-the-art LLMs on transforming textual problem descriptions into executable constraint models. We explored the use of different modelling frameworks, with varying abstraction and interface types. We also explored the use of prompt-based and inference-time compute methods for enhancing the performance of LLMs in constraint modelling.
Our evaluation indicated an advantage of Python-based frameworks (OR-Tools, CPMpy) compared to the domain-specific MiniZinc language, with further improvements when appending detailed guidelines and documentation in system prompts. Reasoning models were shown to be a good fit for this task. Inference-time methods were shown to provide an improvement in general, with repeated sampling and self-verification showing up to a 10% increase in accuracy. Notably, retrieval-augmented in-context learning was not effective when combined with increased documentation in the prompts. The strongest performance came from combining repeated sampling and self-verification, achieving 70% accuracy, suggesting that LLMs can indeed assist in the modelling process, but human expert validation is still required, especially for critical tasks.
The existence of CP-Bench allows for systematically evaluating more design choices, including newer LLMs, additional frameworks (e.g., integer linear programming solvers), more coding LLM techniques (e.g. retrieval-augmented documentation), or supervised fine-tuning; the latter would require a training dataset of constraint models and problem descriptions. While our benchmark contains more diverse combinatorial problems, they often originate from textbooks, whereas industrial problems typically involve more data, complex constraints, and elaborate descriptions—though rarely available publicly. Future work could also explore multi-turn interactions, where a user and system iteratively refine the model. Finally, while this paper focused on evaluating model correctness, experimenting with scaled-up problem instances and solving times is worth investigating, with interesting directions being learning to model efficiently or to re-model for efficiency.
{ack}
This project has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (Grant No. 101002802, CHAT-Opt), and from the Flemish Government under “Onderzoeksprogramma Artificiële Intelligentie (AI) Vlaanderen”.